Katy needs to mix a 30% acid solution with a 60% acid solution to create 200 millileters of a 52.5% solution. How many millileters of each solution must Katy use?
Answer by checkley77(12844)

.60X+.30(200-X)=.525*200
.60+60-.30X=105
.30X=105-60
.30X=45
X=45/.30
X=150 MILLILITERS OF THE 60% ACID SOLUTION IS USED.
PROOF:
.60*150+.30(200-150)=105
90+.30*50=105
90+15=105
105=105